OpenStudy (anonymous):

WHAT ARE THE FACTORS OF THE FOLLOWING TRINOMIAL? s^2 + 10s +25

OpenStudy (anonymous):

What two numbers add to make 10 but multiply to make 25?

OpenStudy (anonymous):

i mean polynomial >?

OpenStudy (anonymous):

No, it's a trinomial.

OpenStudy (anonymous):

You are looking for two numbers which add to make 10 but multiply to make 25.

OpenStudy (anonymous):

5 n 5 and its polynomial cause im looking at the worrk that im doing .

OpenStudy (anonymous):

are they negative or positive sighs

OpenStudy (anonymous):

Good! They are both positive. since otherwise they would not add to 10.

OpenStudy (anonymous):

can yu help me with this one also .4w^2 +20w +25 help me with the steps

OpenStudy (anonymous):

Yes, one moment.

OpenStudy (anonymous):

So we are looking for some value a and b such that (2w + a)(2w + b) = 4w^2 + 20w + 25 That means we need: 2a + 2b = 20 and a * b = 25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5 and 5 also right

OpenStudy (anonymous):

Yep! Nice job.

OpenStudy (anonymous):

3m^2 -12m + 9 help aqain im trying to understand more

OpenStudy (anonymous):

Ok so here we must have a 3 on one of the m terms, and a 1 on the other since no other numbers multiply to make 3. So we need: (3m + a)(m + b) = 3m^2 - 12m + 9 That means that: 3b + a = -12 and a * b = 9

OpenStudy (anonymous):

3 and 3

OpenStudy (anonymous):

12g^2 +42g -24

OpenStudy (anonymous):

i know dis one start with 6g